OpenStudy (gabylovesyou):

The formula for the perimeter P of a rectangle with length l and width w is P = 2l + 2w. Which of the following is a formula for the length of a rectangle in terms of the perimeter and width?

jimthompson5910 (jim_thompson5910):

The idea is to fully isolate L. To do that, you subtract 2w from both sides first, then you divide both sides by 2. \[\large P = 2l + 2w\] \[\large P - 2w = 2l\] \[\large 2l = P - 2w\] \[\large l = \frac{P - 2w}{2}\]